1.Nevada Company manufactures and sells office chairs in two models: Executive and Deluxe. Product information is provided below.

                                                                                     Deluxe        Executive

                        Unit selling price                                   $150               $500

                        Unit variable costs                                      90                  250

                        Unit contribution margin                         $ 60               $250

           

                        Machine hours required per unit                  2                   10

            Nevada has 10,000 machine hours available each month for producing these two products. The demand for both products is such that Nevada can sell as many units of either product as it can produce. What is the maximum amount of contribution margin that the company can obtain from the available machine hours next month?  

  1. $250,000.
  2. $600,000.
  3. $300,000.
  4. $2,500,000.
  1. Referring to the previous problem, what if all of the facts were the same, except there was a limit to customer demand of 3,250 units of the Deluxe and 400 units of Executive. What is the maximum contribution margin that the company can obtain from the available machine hours given the limit to the demand?
  1. $282,500
  2. $300,000
  3. $195,000
  4. $280,000
Answer #1

Calculate contribution margin per hour

Deluxe Executive
Unit contribution margin 60 250
Machine hour per unit 2 10
Contribution margin per machine hour 30 25
Rank 1 2

1) Maximum contribution margin = 10000/2*60 = $300000

So answer is c) $300000

2) 3250 deluxe unit can be made in 3250*2 = 6500 hours

executive unit = (10000-6500)/10 = 350 Unit

Maximum contribution margin = (3250*60+350*250) = $282500

So answer is a) $282500